# Congestion-Aware Path Re-routing Strategy for Dense Urban Airspace

### v1

@@ Abstract
Index Terms:
Uncrewed aircraft system (UAS), UAS Traffic Management (UTM), congestion-aware path planning, Discrete-time queuing theory
I Introduction
@@
With the rapid growth in the number of uncrewed aircraft systems (UAS) being deployed for on-demand delivery applications, severe congestion is foreseen in the urban airspace. There is a need for structuring the airspace to mitigate this congestion. Delivery applications is urban airspaces would normally require UAS to travel on preplanned paths between source-destination pairs. Geofencing these paths can physically separate a UAS trajectory from other UAS, as well as static obstacles. However, intersecting UAS trajectories are inevitable, especially in dense traffic scenarios. Increased traffic on these paths due to increasing delivery demand results in severe conflicts among the UAS. Further, the increasing number of conflicts in confined geofenced air volumes could be detrimental to UAS safety. Such limitations are majorly due to treating UAS traffic as a road-like network or air traffic, where UAS are restricted to the preplanned paths and air volumes. To overcome this problem, we propose in this paper an airspace design and a congestion mitigation strategy that utilizes the airspace surrounding the preplanned path and dynamically generates local paths for UAS. These paths are opted for when the preplanned paths are congested. We also suggest an approach for estimating the expected air volume that would be utilized as a function of the probabilistic occurrence of congestion. Such airspace can be adaptively reserved beforehand to accommodate any foreseen congestion.
@@
Background and Literature: Several government and private agencies [1, 2, 3] have developed different airspace designs to enable beyond-visual range UAS operations. These designs impose structural constraints on UAS’s degree of freedom (DoF) to maximize airspace safety and capacity at the expense of introducing time delays in UAS flights. Broadly, the structures proposed suggest segmenting the airspace and restricting the UAS in an air matrix (discrete UAS DoF) [4]; in a network of virtual air geofenced corridors (unidirectional UAS traffic flow in an enclosed air corridor volume) [5]; or in geo-vectored altitude layers, where UAS heading vector range is limited in each altitude layer and the altitude layers are interconnected by vertical ascend or descend corridors) [6]. UAS safety risk is a probabilistic belief on collision occurrence among UAS. In [7], the event where the separation between UAS is below a threshold is termed an intrusion. An anticipated intrusion in a prescribed look-ahead time window is termed a conflict. Whereas congestion is an accumulation of UAS in a predefined region that eventually leads to increased conflict risk. The airspace designs have an intrinsic ability to reduce the occurrence of collision due to the constraints that it imposes on the UAS traffic.
In air matrix design, the airspace is tessellated into a cellular grid, and online path-planning strategies such as Rapidly exploring Random Trees (RRT), skeleton maps [8, 9] were proposed that generate optimal and risk-aware trajectories for UAS. However, with the increasing number of UAS, these strategies lead to an unstructured and unpredictable traffic pattern.
In altitude-layered airspace, the intrinsic safety can be improved if the UAS is capable of sensing and avoiding (SAA) the intrusions [10] or can mitigate conflict risk by preflight or in-flight conflict detection and resolution (CDR) [11, 12].
@@
Though the SAA and CDR are adequate at individual UAS levels, these approaches perform poorly in congested traffic scenarios. In the study conducted by [13], beyond a critical demand, the SAA can no longer execute avoidance maneuvers without colliding with neighboring UAS. Whereas the in-flight CDR suffers from domino effects and spillback conflicts, which can severely destabilize the UAS traffic [14, 15]. A solution to this problem as proposed in this paper, is to mitigate congestion and prevent high UAS density regions from forming than to allow UAS to proceed and encounter conflicts in dense regions.
@@
The analytical study in [16] suggests segregating UAS into altitude layers as prima facie for mitigating congestion. In [14], when two or more UAS conflict, these UAS are collectively enclosed in a virtual disk and labeled as congested. Any UAS in its vicinity locally re-route, avoiding this disk until the conflicts in the disk are resolved. In [17], the airspace is tesselated into cells, and the cell is said to be congested if the number of pre-planned paths occupying the same cell is greater than the cell capacity. Similarly, in [18], a region is congested if the expectation of UAS encountering conflict with two or more UAS is above some threshold. The papers [19, 20, 21] present strategies for mitigating congestion in the air corridor networks. In [19], the UAS traffic is modeled as fluid queues, and the congestion is mitigated by controlling the UAS inflow and outflow rate for each corridor in the network. In [20], the velocity of the trailing UAS is reduced proportional to the weighted sum of UAS inter-separation distances in the congested volume. In [21], the congestion is mitigated by solving a traffic assignment problem such that the number of UAS routed on each corridor results in an equitable UAS density distribution in all corridors. In [18, 17, 21], the mitigation strategy is spatial and does not capture congestion in the temporal sense, forcing new incoming UAS to take long detours. Whereas in [19, 20], intermittent local congestion events affect the throughput of the entire corridor network. This limitation is because the UAS is always restricted within air corridors despite airspace being available in the surrounding airspace.
@@
The paper is organized as follows. The workspace for the congestion mitigation problem is defined in Section II. Section III introduces congestion, the heading structural constraints being imposed, and the proposed congestion mitigation strategy employed by UAS. Section IV presents queuing models to study the emerging traffic behavior as a function of the UAS traffic demand and structural constraints. Section V presents simulations that demonstrate the increased intrinsic safety and resemblance of UAS traffic to the parallel air corridor network. Section VI concludes the paper, emphasizing the advantages of the proposed mitigation strategy.
II Problem Definition
@@
This paper presents traffic analysis for a UAS traffic management (UTM) framework driven by stochastic delivery demand. The UTM environment has static obstacles and several source-destination pairs. The delivery service provider deploys a UAS for every customer requesting delivery on a specific source-destination route, and such requests recur stochastically. Accordingly, at the source location, a time-varying geometrically distributed request for UAS is assumed with bounded request rate $\Lambda(t)\in[0,\Lambda_{\max}]$, where $\Lambda_{\max}$ is the maximum request rate. The time axis is divided into slots of $\Delta T=1/\Lambda_{\max}$ time units and indexed with $k\in\mathbb{N}_{0}$. We assume that though multiple requests are received in a given time slot, only one request among them is randomly chosen, for which a UAS is deployed at the start of the next timeslot. The remaining requests are dropped. The probability that a UAS would be deployed in a timeslot is given by $\lambda=\Lambda\Delta T$.
@@
The workspace is assumed to be a bounded horizontal stretch of air volume containing the given source-destination pair. It is tessellated into a connected grid of a finite number of identical cells. As the static obstacles are of varying heights, some may intersect the cells of this workspace. We define a geometric center called the cell center for each static obstacle-free cell (refer Fig. 1). The motion of a UAS leaving a cell center and traveling in a straight line to reach a neighboring cell center is referred to as a transition. The traversal of UAS in the workspace is modeled as successive transitions from one cell center to any of its neighboring cell centers. We assume that the UAS takes $\Delta T$ seconds to transition between neighboring cell centers. Two UAS that are $\tau$ cells spatial separated are equivalently $\tau$ slots ($\tau\Delta T$) temporal separated.
Figure 1: The cell tessellation of workspace containing the UAS dispatch location (source) and the payload delivery location (destination).
The UTM is assumed to prescribe a nominal path connecting cell centers between any source-destination pair. We assume this path would be approximated by connected straight-line segments passing through cell centers. Since the UAS requests are assumed geometrically distributed with request rate $\lambda$, the inter-separation distance between UAS transitioning on the nominal path (in time slots) is also geometrically distributed and is given by the distribution $A(t)$ as follows.
@@
The requests are recurring, hence there is a stream of UAS traffic flowing in the nominal path from source to destination. The traffic in the path, right ahead in front of a UAS, is its upstream traffic. The traffic trailing behind the UAS is its downstream traffic. When $\lambda$ is large, $\mathrm{P}[\tau\leq S]$ is high; thus, the chances of a UAS encountering congestion on its upstream is high. When the probability of congestion increases, there is a requirement to generate alternative parallel paths (dependent on the request rate) that avoid congested regions.
@@
The nominal path is a reference path given to the UAS. However, the UAS is free to opt for any of those cell centers in its neighborhood that do not violate the constraints set by the UTM. The main objective of this paper is to propose a distributed congestion mitigation strategy for the UAS. Using this strategy, individual UAS can dynamically generate local parallel paths for itself in the unoccupied airspace in a structured manner, following which the UAS bypasses congested regions in its local neighborhood. When the request rate $\lambda$ is low, the UAS traverses along the nominal flight path. However, in the event of congestion on the nominal path, a certain number of UAS present in the downstream are diverted onto the parallel paths. As the request rate increases, the frequency of the congestion event increases. The number of UAS that would be diverted on parallel paths also increases, causing congestion on the parallel paths as well. Then, UAS downstream would have to opt for parallel paths further outward from the nominal path. When congestion reduces, the UAS on parallel paths tends to close in on the nominal path. In this process of mitigating congestion, there would be a change in the area occupied by the UAS stream around the nominal path with variation in the request rate.
@@
The above idea is illustrated in Fig. 2. The UAS only relies on the nominal path information from UTM and the positional information of UAS present in its vicinity to determine congested regions and generate local parallel paths. For the UAS traffic streams routed between source-destination pairs, we intend to demonstrate and analyze the extent of the UAS traffic spread around the nominal path as a function of the request rate. The paper also extends congestion mitigation to scenarios where the inter-separation distances between UAS of intersecting UAS traffic streams fall below the threshold $S$.
@@
Figure 2: Several feasible paths may exist between source and destination. Here, the path through 1-4 is the nominal path suggested by UTM among the feasible paths. In the event of congestion on this path, the figure illustrates the local parallel paths that are generated. The average number of parallel paths on which the UAS traffic would be diverted depending on the request rate $\lambda$ is the expected traffic spread.
III Congestion Mitigation Methodology
@@
The nominal path between the source-destination pair is segmented into a connected chain of straight-line segments. The set of cells through which the straight-line segment passes is called a nominal segment, with a start and an end cell defined for each segment. The proposed strategy deals with congestion in each nominal segment separately. A unit vector directed from start to end is called an upstream vector, and that directed from end to start is called a downstream vector. To identify a region as congested, we first need to geometrically define the region, for which, we partition the workspace as follows.
@@
Assume the UAS are identical in all aspects and have a $L$ slot look-ahead time window to predict where other UAS would be in the next $L\Delta T$ time ahead. The cells in the nominal segment (start and end included) that are $S=2L+1$ time slots apart are called nodes (shown in Fig. 3(a)). With each node as a geometric center, the set of cells enclosed in a square region with edge length $S$ timeslots (an edge parallel to the segment) is called a nominal zone. The eight square regions (of $S$ timeslots edge length) surrounding the nominal zone are called zones and form the neighborhood of the nominal zone. Similarly, we may define the neighborhood for any arbitrary zone. We assume a finite grid of identical zones for each nominal segment to facilitate alternative path generation in the event of congestion. The grid is so positioned that the nominal zones are a subset of the grid and bisect the grid (shown in Fig. 3(b)). The length of the grid (that is the length of the nominal segment) is significantly larger than the width. The start cell, end cell, and width of the grid for two connected nominal segments are chosen such that the respective grids do not intersect. A stream-level tuple notation $Z=(X,Y)$ is used for referring to zones in the finite grid, where Stream $X\in\{-X_{e},...,0,...,X_{e}\}$ and Level $Y\in\{1,...,Y_{e}\}$ and, $X_{e}$ and $Y_{e}$ are positive integers (refer Fig. 3(b)). The nominal zones are called Stream 00 zones denoted with tuple $(0,Y)$, where $Y$ denotes the zone level in the upstream direction. $(0,1)$, $(0,Y_{e})$ denote the nominal zones containing start and end cells, respectively. We define the congested regions as follows.
@@
Figure 3: a) Node and zone definition along the flight path. b) Connected grid and stream-level indexing of zones about each nominal segment
@@
A zone is said to be congested if $M$ or more than $M$ number of UAS are present in the zone.
$L$ and $M$ are UTM design parameters. Note that the congestion in the zone is due to UAS transitioning on the given source-destination route as well as due to any exogenous UAS coexisting in the workspace. The grid construction is UAS-agnostic. The UAS can extract the grid geometry from the nominal segment information shared by the UTM. Within the zone, the heading of the UAS would be restricted so that UAS would only transition on a subset of cells. The heading constraint is position-dependent, for example, when UAS is in the nominal segment it is restricted to transition in the upstream direction. A UAS in a zone can detect or communicate with other UAS coexisting in the zone and its respective neighborhood. By knowing the UAS present positions, the heading constraint imposed, and that any UAS would complete one transition in one timeslot, we assume the UAS using a linear kinematic model can predict other UAS state information (position and heading) $L\Delta T$ time ahead with sufficient accuracy.
The UAS on reaching a node needs $L$ transitions on the nominal segment to reach the neighboring zone. Since a $L$ slot look-ahead time window is assumed, when the UAS is at a node and finds the neighboring zone is likely to turn congested in the next $L$ slots, then the UAS could decide to avoid the neighboring zone.
@@
From the perspective of a UAS in zone $Z$ (depicted in Fig. 4), the neighboring zone $Z^{\prime}$ will likely be congested if the number of $L$ slot look-ahead UAS predicted positions in zone $Z^{\prime}$ is greater than or equal to $M$.

### v2

# Conceptualization of this study, Methodology, Software, Formal Analysis, Writing-Original Draft

@@ Abstract
keywords:
[…]
1 Introduction
@@
With the rapid growth in the number of uncrewed aircraft systems (UAS) being deployed for on-demand delivery applications, severe congestion is foreseen in the urban airspace. There is a need for structuring the airspace to mitigate this congestion. Delivery applications in urban airspaces would normally require UAS to travel on preplanned paths between source-destination pairs. Geofencing these paths can physically separate a UAS trajectory from other UAS, as well as static obstacles. However, intersecting UAS trajectories are inevitable, especially in dense traffic scenarios. Increased traffic on these paths due to increasing delivery demand results in severe conflicts among the UAS. Further, the increasing number of conflicts in confined geofenced air volumes could be detrimental to UAS safety. Such limitations are majorly due to treating UAS traffic as a road-like network or air traffic, where UAS are restricted to the preplanned paths and air volumes. To overcome this problem, we propose in this paper an airspace design and a congestion mitigation strategy that utilizes the airspace surrounding the preplanned path and dynamically generates local paths for UAS. These paths are opted for when the preplanned paths are congested. We also suggest an approach for estimating the expected air volume that would be utilized as a function of the probabilistic occurrence of congestion. Such airspace can be adaptively reserved beforehand to accommodate any foreseen congestion.
@@
Background and Literature: Several government and private agencies (Kopardekar et al. (2016); Catapult (2020); Balakrishnan et al. (2018)) have developed different airspace designs to enable beyond-visual range UAS operations. These designs impose structural constraints on UAS’s degree of freedom (DoF) to maximize airspace safety and capacity at the expense of introducing time delays in UAS flights. Broadly, the structures proposed suggest segmenting the airspace and restricting the UAS in an air matrix (discrete UAS DoF) (Pang et al. (2020)); in a network of virtual air geofenced corridors (unidirectional UAS traffic flow in an enclosed air corridor volume) (Tony et al. (2021)); or in geo-vectored altitude layers, where UAS heading vector range is limited in each altitude layer and the altitude layers are interconnected by vertical ascend or descend corridors) (Hoekstra et al. (2018)).
UAS safety risk is a probabilistic belief on collision occurrence among UAS. In Doole et al. (2021), the event where the separation between UAS is below a threshold is termed an intrusion. An anticipated intrusion in a prescribed look-ahead time window is termed a conflict. Whereas congestion is an accumulation of UAS in a predefined region that eventually leads to increased conflict risk. The airspace designs have an intrinsic ability to reduce the occurrence of collision due to the constraints that it imposes on the UAS traffic.
@@
In air matrix design, the airspace is tessellated into a cellular grid, and online path-planning strategies such as Rapidly exploring Random Trees (RRT), skeleton maps (Li et al. (2023); Abdul et al. (2023)) were proposed that generate optimal and risk-aware trajectories for UAS. However, with the increasing number of UAS, these strategies lead to an unstructured and unpredictable traffic pattern. In altitude-layered airspace, the intrinsic safety can be improved if the UAS is capable of sensing and avoiding (SAA) the intrusions (Agnel Tony et al. (2021)) or can mitigate conflict risk by preflight or in-flight conflict detection and resolution (CDR) (Sacharny et al. (2022); Nagrare et al. (2024)). Though the SAA and CDR are adequate at individual UAS levels, these approaches perform poorly in congested traffic scenarios. In the study conducted by Watkins et al. (2021), beyond a critical demand, the SAA can no longer execute avoidance maneuvers without colliding with neighboring UAS. Whereas the in-flight CDR suffers from domino effects and spillback conflicts, which can severely destabilize the UAS traffic (Sedov et al. (2017); Aarts et al. (2023)). A solution to this problem as proposed in this paper, is to mitigate congestion and prevent high UAS density regions from forming than to allow UAS to proceed and encounter conflicts in dense regions.
@@
The analytical study in Sunil et al. (2018) suggests segregating UAS into altitude layers as prima facie for mitigating congestion. In Sedov et al. (2017), when two or more UAS conflict, these UAS are collectively enclosed in a virtual disk and labeled as congested. Any UAS in its vicinity locally re-route, avoiding this disk until the conflicts in the disk are resolved. In Lee et al. (2022), the airspace is tesselated into cells, and the cell is said to be congested if the number of pre-planned paths occupying the same cell is greater than the cell capacity. Similarly, in Egorov et al. (2019), a region is congested if the expectation of UAS encountering conflict with two or more UAS is above some threshold. The papers Zhou et al. (2020); Gharibi et al. (2021); Wang et al. (2021) present strategies for mitigating congestion in the air corridor networks. In Zhou et al. (2020), the UAS traffic is modeled as fluid queues, and the congestion is mitigated by controlling the UAS inflow and outflow rate for each corridor in the network. In Gharibi et al. (2021), the velocity of the trailing UAS is reduced proportional to the weighted sum of UAS inter-separation distances in the congested volume. In Wang et al. (2021), the congestion is mitigated by solving a traffic assignment problem such that the number of UAS routed on each corridor results in an equitable UAS density distribution in all corridors. In Egorov et al. (2019); Lee et al. (2022); Wang et al. (2021), the mitigation strategy is spatial and does not capture congestion in the temporal sense, forcing new incoming UAS to take long detours. Whereas in Zhou et al. (2020); Gharibi et al. (2021), intermittent local congestion events affect the throughput of the entire corridor network. This limitation is because the UAS is always restricted within air corridors despite airspace being available in the surroundings.
@@
The paper is organized as follows. The workspace for the congestion mitigation problem is defined in Section 2. Section 3 introduces congestion, the heading structural constraints being imposed, and the proposed congestion mitigation strategy employed by UAS. Section 4 presents queuing models to study the emerging traffic behavior as a function of the UAS traffic demand and structural constraints. Section 5 presents simulations that demonstrate the increased intrinsic safety and resemblance of UAS traffic to the parallel air corridor network. Section 6 concludes the paper, emphasizing the advantages of the proposed mitigation strategy.
2 Problem Definition
@@
This paper presents traffic analysis for a UAS traffic management (UTM) framework driven by stochastic delivery demand. The UTM environment has static obstacles and several source-destination pairs. The delivery service provider deploys a UAS for every customer requesting delivery on a specific source-destination route, and such requests recur stochastically. Accordingly, at the source location, a time-varying geometrically distributed request for UAS is assumed with bounded request rate $\Lambda(t)\in[0,\Lambda_{\max}]$, where $\Lambda_{\max}$ is the maximum request rate. The time axis is divided into slots of $\Delta T=1/\Lambda_{\max}$ time units and indexed with $k\in\mathbb{N}_{0}$. We assume that though multiple requests may be received in a given timeslot, only one request among them is randomly chosen, for which a UAS is deployed at the start of the next timeslot. The probability that a UAS would be deployed in a timeslot is given by $\lambda=\Lambda\Delta T$.
@@
The workspace is assumed to be a bounded horizontal stretch of air volume containing the given source-destination pair. It is tessellated into a connected grid of a finite number of identical cells. As the static obstacles are of varying heights, some may intersect the cells of this workspace. We define a geometric center called the cell center for each static obstacle-free cell (refer Fig. 1(a)). The motion of a UAS leaving a cell center and traveling in a straight line to reach a neighboring cell center is referred to as a transition. The traversal of UAS in the workspace is modeled as successive transitions from one cell center to any of its neighboring cell centers. We assume that the UAS takes $\Delta T$ time units to transition between neighboring cell centers. Two UAS that are $\tau$ cells spatial separated are equivalently $\tau$ slots ($\tau\Delta T$) temporal separated. The UTM prescribes a nominal path connecting cell centers between any source-destination pair. We assume this path would be approximated by connected straight-line segments passing through cell centers. Since the UAS deployment is geometrically distributed with parameter $\lambda$, the inter-separation distance between UAS transitioning on the nominal path (in timeslots) is also geometrically distributed and is given by the distribution $A(t)$.
@@
The requests are recurring, hence there is a stream of UAS traffic flowing in the nominal path from source to destination. The traffic in the path, right ahead in front of a UAS, is its upstream traffic. The traffic trailing behind the UAS is its downstream traffic. When $\lambda$ is large, $\mathrm{P}[\tau\leq S]$ is high; thus, the chances of a UAS encountering congestion on its upstream is high. When the probability of congestion increases, there is a requirement to generate alternative parallel paths (dependent on the request rate) that avoid congested regions. The nominal path is a reference path given to the UAS. However, the UAS is free to opt for any of those cell centers in its neighborhood so long as the choice does not violate the constraints set by the UTM. The main objective of this paper is to propose a distributed congestion mitigation strategy for the UAS. Using this strategy, individual UAS can dynamically generate local parallel paths for itself in the unoccupied airspace in a structured manner, following which the UAS bypasses congested regions in its local neighborhood. When the request rate $\lambda$ is low, the UAS traverses along the nominal flight path. However, in the event of congestion on the nominal path, a certain number of UAS present in the downstream are diverted onto the parallel paths. As the request rate increases, the frequency of the congestion event increases. The number of UAS that would be diverted on parallel paths also increases, causing congestion on the parallel paths as well. Then, UAS downstream would have to opt for parallel paths further outward from the nominal path. When congestion reduces, the UAS on parallel paths tends to close in on the nominal path. As the request rate is time-varying, a change in the area occupied by the UAS stream around the nominal path will be observed in the process of congestion mitigation.
@@
The above idea is illustrated in Fig. 1(b). The UAS only relies on the nominal path information from UTM and the positional information of UAS present in its vicinity to determine congested regions and generate local parallel paths. For the UAS traffic routed between source-destination pairs, we demonstrate and analyze the extent of the UAS traffic spread around the nominal path as a function of the request rate. The paper extends congestion mitigation to scenarios where the inter-separation distances between $M$ UAS of intersecting traffic streams fall below the threshold $S$.
@@
Figure 1: a) Cell tessellation of workspace containing the UAS dispatch center (source) and the customer location (destination). b) Several feasible paths may exist between source and destination ($\mathsf{S}$-$\mathsf{1}$-$\mathsf{2}$-$\mathsf{4}$-$\mathsf{D}$, $\mathsf{S}$-$\mathsf{1}$-$\mathsf{3}$-$\mathsf{4}$-$\mathsf{D}$, $\mathsf{S}$-$\mathsf{5}$-$\mathsf{4}$-$\mathsf{D}$). Here, the path shown by the bold line is the nominal path given by the UTM. In the event of congestion on this path, the figure illustrates the local parallel paths that are generated. The expected traffic spread or the expected path spread is the average number of parallel paths on which the UAS traffic would be diverted depending on the request rate $\Lambda$.
3 Congestion Mitigation Methodology
@@
The nominal path between the source-destination pair is segmented into a connected chain of straight-line segments. The set of cells through which the straight-line segment passes is called a nominal segment, with a start and an end cell defined for each segment. The proposed strategy deals with congestion in each nominal segment separately. A unit vector directed from start to end is called an upstream vector, and that directed from end to start is called a downstream vector. To identify a region as congested, we first need to geometrically define the region, for which we partition the workspace as follows.
@@
Assume the UAS are identical in all aspects and have a $L$ slot look-ahead time window to predict where other UAS would be in the next $L\Delta T$ time ahead. The cells in the nominal segment (start and end included) that are $S=2L+1$ timeslots apart are called nodes (shown in Fig. 2a). With each node as a geometric center, the set of cells enclosed in a square region with edge length $S$ timeslots (an edge parallel to the segment) is called a zone. The eight square regions (of $S$ timeslots edge length) surrounding the zone are called neighborhood zones. We may successively define the neighborhood for all neighborhood zones. The zone and its corresponding node are denoted with $Z$ and $\widetilde{Z}$, respectively. We assume a finite grid of identical zones for each nominal segment to facilitate alternative path generation in the event of congestion. The grid is so positioned that the nominal segment bisects the grid (shown in Fig. 2). The length of the grid (that is, the length of the nominal segment) is significantly larger than the width. The start cell, end cell, and width of the grid for two connected nominal segments are chosen such that the respective grids do not intersect. The zones through which the nominal segment passes are called nominal zones. A stream-level tuple notation $Z=(X,Y)$ is used for referring to zones in the finite grid, where stream $X$ is abscissa, $X\in\{-X_{e},...,0,...,X_{e}\}$ and the level $Y$ is ordinate, $Y\in\{1,...,Y_{e}\}$, and $X_{e}$ and $Y_{e}$ are positive integers (refer Fig. 2b top view). The nominal zones are called Stream$(0)$ zones denoted with tuple $(0,Y)$, where $Y\in\{1,...,Y_{e}\}$ is the zone level in the upstream direction. Here, $(0,1)$, $(0,Y_{e})$ are the nominal zones containing start and end cells, respectively. Similarly, Stream(±$(\pm 1)$, Stream(±$(\pm 2)$, and so on Stream(±$(\pm X_{e})$ zones can be identified. By interconnecting nodes of the Stream$(X)$ zones with respective upstream nodes, $X\in\{-X_{e},...,X_{e}\}/\{0\}$, we get $2X_{e}$ parallel segments to the nominal segment, referred to as Stream$[X]$ segments. The congested regions are defined as follows.
@@
Figure 2: a) Node and zone definition. b) Connected grid and stream-level indexing of zones about each nominal segment.
@@
A zone is said to be congested if $M$ or more than $M$ UAS are present in the zone.
Here, $L$ and $M$ are UTM design parameters. Note that the congestion in the zone is due to UAS transitioning on the given source-destination route as well as due to any exogenous UAS coexisting in the workspace. The grid construction is UAS-agnostic. The UAS can extract the grid geometry from the nominal segment information and design parameters shared by the UTM. Within the zone, the heading of the UAS would be restricted so that UAS would only transition on a subset of cells. The heading constraint is position-dependent. For example, when UAS is in the nominal segment, it is restricted to transition in the upstream direction. A UAS in a zone can detect or communicate with other UAS coexisting in the zone and respective zone neighborhood. By knowing the UAS present positions, the heading constraint imposed, and that any UAS would complete one transition in one timeslot, we assume the UAS using a linear kinematic model can predict other UAS state information (position and heading) $L\Delta T$ time ahead with sufficient accuracy.
The UAS on reaching a node, needs to take at least $L$ successive transitions to enter any of the neighboring zones. The UAS is equipped with $L$ slot look-ahead prediction window. The UAS can predict other UAS positions and determine which neighboring zones would probably become congested in the next $L$ slots. Thus, when the UAS is at the node, it could selectively choose not to transition towards neighboring zones that will become congested.
@@
From the perspective of a UAS in zone $Z$ (depicted in Fig. 3), the neighboring zone $Z^{\prime}$ is congested if the number of $L$ slot look-ahead UAS predicted positions in zone $Z^{\prime}$ is greater than or equal to $M$.